# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ I.1 Bounds on Number of Floppy Modes
For the second bound, we start with a lattice of $N_{2}=0$ and all triangles as $T_{1}$; In this limit, the maximal number of loops givs $L\leq\frac{N}{6}$, by arranging the loops in a honeycomb formation. Now, as we gradually switch $T_{1}$ blocks to $T_{2}$, the most efficient way to create new loops is that two loops share each $T_{2}$ block. Each new loop needs six bonds, thus $L\leq\frac{N}{6}+\frac{N_{2}}{3}$, and $F\leq\frac{2N_{1}}{3}+\frac{P}{2}$. The crossover between these two bounds occurs at $N_{1}=\frac{3P}{2}-3$, and we obtain the upper bound given in Eq. (2).
For a given lattice of size and shape, the system becomes more rigid as there are more $T_{2}$ than $T_{1}$ blocks. We obtain the minimal number $N_{2}$ of $T_{2}$ blocks or maximal number $N_{1}$ of $T_{1}$ blocks required for rigidity by requiring that all $E$ edge nodes belong to a single chain with at least one odd loop. Thus, the number of internal bonds must be at least $E$, with $E-1$ bonds connecting all nodes and one bond closing an odd loop. Thus, mechanical rigidity can be achieved when $N_{1}\leq\frac{N-P}{2}$. In this case $F=0$. As we increase $N_{1}$, we remove internal bonds. There are no loops, and this removal of bonds does not create loops. Thus, by substituting in Eq. (1) we obtain the lower bound given in Eq. (2).
@@
and the upper bound by $F\leq\frac{2N_{1}}{3}$. As system size increases, the average number of floppy modes in randomly generated lattices collapses to a single curve, see Fig. 5.
Figure 5: Average number of floppy modes in randomly generated lattices. a, Results for different system sizes normalized by $\tilde{F}=F/[(N+P)/2]$ and $\tilde{N_{1}}=[N_{1}-(N-P)/2]/[(N+P)/2]$, such that the lower bound will coincide. b, Normalized umber of floppy modes at the onset of rigidity, $\tilde{N}_{1}=0$.
I.2 Matrix-Vector Multiplication with Multi-Layer Rigid Loops
We have demonstrated the matrix-vector multiplication in a single-layer metamaterial. Here, we further show a larger matrix-vector multiplication in a three-layer metamaterial of four inputs in the middle layer and 12 outputs in the side layers. These frustrated loops similarly relate the input and output vectors. We experimentally apply a single input and measure the transmission matrix $A$ from the slopes of the input-output curves (Fig. 6). Similarly to the single-layer case, we predict the transmission by assuming decay factor $\alpha=0.8$ within each layer, decay factor $\beta_{1}=0.4$ when the input transmits from one layer to another layer, and $\beta_{2}=1$ when two layers have the same input, like the output in the left layer $O_{\rm T}=I_{\rm T}$ (Fig. 4d). We find good agreement between the theoretical predictions and the experimental measurements (Fig. 6).
@@ I.2 Matrix-Vector Multiplication with Multi-Layer Rigid Loops
Figure 7: 3D printed metamaterials and experimental method. a, A 3D printed metamaterial for buckling with elastoplastic material (Nylon). b, A 3D printed metamaterial for buckling with rubber-like polymer (Polyject Agilus + Vero). c, A metamaterial with tracking markers at the connected joints. d, A single-layer metamaterial for local actuation is printed with two materials. Purple ligaments and red diamonds are printed with soft material (Agilus) and rigid material (Vero), respectively. e, Experimental setup of a three-layer metamaterial with local actuation under filming of two cameras. The picture in the top-right corner is the recorded image from the right camera
@@ I.3 Experimental methods
The metamaterials with floppy modes designed in this work consist of a plurality of triangular unit cells, wherein the rigid bonds and rotating nodes from the theoretical model are replaced with solid diamond blocks of length $\ell$ and ligaments of thickness $t$, respectively. For the metamaterials with buckling floppy modes, we 3D printed them with a single material of $\ell=10.3$ mm, $t=0.4$ mm, and $d=3$ mm. For the metamaterials with elastoplastic buckling (Fig. 7a), we used an elastoplastic material, nylon, with the fused deposition modelling (FDM) technology (Ultimaker 2S). For the metamaterials with elastic buckling (Fig. 7b), we 3D printed them with a rubber-like polymer material (Stratasys, Aligus+Vero). The thicknesses of the metamaterials are set to $T=24.5$ mm for elastoplastic metamaterials and $T=23.2$ mm for elastic metamaterials to prevent out-of-plane buckling. We then marked the ligaments on one side of the metamaterials with white dots for image tracking (Fig. 7c). We used Image-J software to track the position of the white dots of each frame in the video. For the metamaterials with local actuation, we changed the design of ligaments from a rectangular section to a section with inserted tapered rigid parts (Fig. 7d). We then printed such ligaments with a soft material (Aligus) and the diamonds with a rigid material (Vero). Such ligament design can significantly increase resistance to shear and tension while offering only a negligible resistance to bending  Bossart et al. (2021). For the single-layer metamaterial (Fig. 4b), we set the thickness of the sample to 6 mm. For the three-layer metamaterials with multiple inputs and outputs, we printed each layer separately and assembled them with two rigid frames. The middle layer has a thickness of 6 mm and the left and right layers have a thickness of 3.6 mm. We set a small distance of 0.5 mm between each layer to avoid friction during actuation. We then fix the sample on the table. We marked the ligaments at the six edges of the hexagon layer with black dots (Fig. 7e top-right corner). We then manually pull and push the rigid bars of the middle layer and record the movement of the 12 dots from the left and right layers with two Basler cameras (Fig. 7e).

### v2

@@ I.1 Bounds on Number of Floppy Modes
Figure 5: Average number of floppy modes in randomly generated lattices. a, Results for different system sizes normalized as $\tilde{F}=F/[(N+P)/2]$ and $\tilde{N_{1}}=[N_{1}-(N-P)/2]/[(N+P)/2]$, such that the lower bound will coincide. b, Normalized umber of floppy modes at the onset of rigidity, $\tilde{N}_{1}=0$.
@@
For the second bound, we start with a lattice of $N_{2}=0$ and all triangles as $T_{1}$; In this limit, the maximal number of loops gives $L\leq\frac{N}{6}$, by arranging the loops in a honeycomb formation. Now, as we gradually switch $T_{1}$ blocks to $T_{2}$, the most efficient way to create new loops is that two loops share each $T_{2}$ block. Each new loop needs six bonds, thus $L\leq\frac{N}{6}+\frac{N_{2}}{3}$, and $F\leq\frac{2N_{1}}{3}+\frac{P}{2}$. The crossover between these two bounds occurs at $N_{1}=\frac{3P}{2}-3$, and we obtain the upper bound given in Eq. (1).
For a given lattice of size and shape, the system becomes more rigid as there are more $T_{2}$ than $T_{1}$ blocks. We obtain the minimal number $N_{2}$ of $T_{2}$ blocks or maximal number $N_{1}$ of $T_{1}$ blocks required for rigidity by requiring that all $E$ edge nodes belong to a single chain with at least one odd loop. Thus, the number of internal bonds must be at least $E$, with $E-1$ bonds connecting all nodes and one bond closing an odd loop. Thus, mechanical rigidity can be achieved when $N_{1}\leq\frac{N-P}{2}$. In this case $F=0$. As we increase $N_{1}$, we remove internal bonds. There are no loops, and this removal of bonds does not create loops. Thus, by substituting in the expression for the number of modes, we obtain the lower bound given in Eq. (1).
@@
and the upper bound by $F\leq\frac{2N_{1}}{3}$. As system size increases, the average number of floppy modes in randomly generated lattices collapses to a single curve (Fig. 5).
I.2 Multi-Layer Matrix-Vector Multiplication
We have demonstrated the matrix-vector multiplication in a single-layer metamaterial. Here, we further show a larger matrix-vector multiplication in a three-layer metamaterial of four inputs in the middle layer and 12 outputs in the side layers. These frustrated loops similarly relate the input and output vectors. We experimentally apply a single input and measure the transmission matrix $A$ from the slopes of the input-output curves (Fig. 6). Similarly to the single-layer case, we predict the transmission by assuming decay factor $\alpha=0.8$ within each layer, decay factor $\beta_{1}=0.4$ when the input transmits from one layer to another layer, and $\beta_{2}=1$ when two layers have input and output at the same location, like the output $L_{\rm T}$ for the input $I_{\rm T}$ (Fig. 4d). We find good agreement between the theoretical predictions and the experimental measurements (Fig. 6).
@@ I.2 Multi-Layer Matrix-Vector Multiplication
Figure 7: 3D printed metamaterials and experimental method. a, Metamaterial for buckling with elastoplastic material (Nylon). b, Metamaterial with tracking markers at the connected joints. c, Stretching stress-strain curve of 3D printed Nylon where we linear fit the elastic modulus and tangent modulus. The white stickers on the dogbone sample are used to track the strain. d, A single-layer metamaterial for local actuation is printed with two materials. Purple ligaments and red diamonds are printed with soft material (Agilus) and rigid material (Vero), respectively. e, Experimental setup of a three-layer metamaterial with local actuation under filming of two cameras. The picture in the top-right corner is the recorded image from the right camera
@@ I.3 Experimental Methods
The metamaterials with floppy modes designed in this work consist of a plurality of triangular unit cells, wherein the rigid bonds and rotating nodes from the theoretical model are replaced with solid diamond blocks of length $\ell$ and width $d$ and ligaments of thickness $t$, respectively. For the metamaterials with buckling floppy modes, we 3D printed them with a single material of $\ell=10.3$ mm, $t=0.4$ mm, and $d=3$ mm (Fig. 7a). For the metamaterials with yield buckling, we used an elastoplastic material, nylon (with a low ratio of $E_{t}/E=0.56\%$ see calibration in Fig. 7c), with the fused deposition modelling (FDM) technology (Ultimaker 2S). For the metamaterials with elastic buckling, we 3D printed them with a rubber-like polymer material (Stratasys, Agilus+Vero). The thicknesses of the metamaterials are set to $D=24.5$ mm for elastoplastic metamaterials and $D=23.2$ mm for elastic metamaterials to prevent out-of-plane buckling. We then marked the ligaments on one side of the metamaterials with white dots for image tracking (Fig. 7b). We used Image-J software to track the position of the white dots of each frame in the video. For the metamaterials with local actuation, we changed the design of ligaments from a rectangular section to a section with inserted tapered rigid parts (Fig. 7d). We then printed such ligaments with a soft material (Agilus) and the diamonds with a rigid material (Vero). Such ligament design can significantly increase resistance to shear and tension while offering only a negligible resistance to bending [39]. For the single-layer metamaterial (Fig. 4b), we set the thickness of the sample to 6 mm. For the three-layer metamaterials with multiple inputs and outputs, we printed each layer separately and assembled them with two rigid frames. The middle layer has a thickness of 6 mm and the left and right layers have a thickness of 3.6 mm. We set a small distance of 0.5 mm between each layer to avoid friction during actuation. We then fixed the sample on the table. We marked the ligaments at the six edges of the hexagon layer with black dots (Fig. 7e top-right corner). We then manually pulled and pushed the rigid bars of the middle layer and recorded the movement of the 12 dots from the left and right layers with two Basler cameras (Fig. 7e).